OpenStudy (anonymous):

the sum of two numbers is 30 and their difference is 6.Find the number.Let x be the large number making x-? the smaller number

OpenStudy (bibby):

So you know that x+(x-?)=30 x-(x-?) = 6 You now have 2 equations

OpenStudy (bibby):

For easier reading we'll pretend x-? is the same thing as y x+y=30 x-y=6->x=6+y 6+y+y=30 6+2y =30 2y =24 y = x-? = 12